OpenStudy (anonymous):

Solve for x x= log base e radical e x=_____a^0

OpenStudy (zehanz):

I think this is what you mean:\(x=\ln \sqrt{e}\)

OpenStudy (anonymous):

x +(log base e) radical 3

OpenStudy (zehanz):

To solve it, write \(\sqrt{e}=e^{\frac{1}{2}}\)

OpenStudy (anonymous):

x = (log base e) radical 3

OpenStudy (anonymous):

this is what i mwant x = log (base e) radical 3

OpenStudy (anonymous):

meant*

OpenStudy (zehanz):

OK, "log base e" is ln, so you have: \(x=\ln \sqrt{3} \Leftrightarrow x=\ln 3^{\frac{1}{2}}\)

OpenStudy (zehanz):

There is a rule for logarithms that says: \(\log_{a}b^k=k \cdot \log_{a}b\). You can use it here.

OpenStudy (zehanz):

So \(x=\dfrac{1}{2} \ln3\)

OpenStudy (zehanz):

Or, if you wish: \(x=\dfrac{1}{2}\log_{e}3\)

OpenStudy (anonymous):

thank you for that formula thnks

OpenStudy (zehanz):

YW!